# Colloquium: Synthetic quantum matter in non-standard geometries

### v1

@@ III.2.1 Charge-pumping, Chern numbers, and gap labelling in 1D quasicrystals
In the case of 1D quasicrystals, the physical relevance of non-zero Chern numbers is evidenced by pumping experiments, where a quantized amount of “charge” is pumped through the chain when the cyclic Hamiltonian parameter is adiabatically changed. The amount of the pumped “charge” is defined by the Chern numbers of the filled bands. Such topological charge pumping has already been proposed by Thouless [243] in the context of 1D periodic (super)lattices. It has been realized in different photonic quasicrystals built from coupled waveguides in Kraus et al. [131], Verbin et al. [254] by adiabatically sweeping the phase of the quasiperiodic potential along the propagation axis. This setup has also served to visualize the presence or absence of localized states at the boundary between different quasicrystals [254], depending on whether the quasicrystals are topologically equivalent or not, see Fig. 10(a). With cold atoms, one can fill the lowest band either as a fermionic band insulator [169] or a bosonic Mott insulator [149] and obtain a quantized charge transport of one site per modulation cycle.
@@ III.2.4 Topological edge states in fractals and quasicrystals
New topological phases are also expected or already observed in fractal and hyperbolic lattices. Whereas for standard geometries, the bulk-edge correspondence establishes an important relation between topological bulk behavior (Chern numbers) and topological edge states, extended bulk regions are lacking in fractal systems. Non-interacting topological systems (i.e., topological insulators and superconductors) on Sierpiński fractals have been studied theoretically in Brzezińska et al. [173], Pai and Prem [183], Iliasov et al. [112], Fremling et al. [86]. In these systems, the Hall conductivity is neither quantized nor proportional to the Chern number of the filled bands. However, splitting into various sub-bands makes it difficult to discern whether a band is filled. The fate of edge states is particularly intriguing since, on a fractal lattice, most sites are near some edge of the system. Topological edge states in a Sierpiński gasket have recently been observed in a photonic setup [30, 143], see Fig. 10(b).
@@
Theoretically simpler than the FQH effect are topological superconductors, such as the paradigmatic Kitaev chain [123]. Even on a mean-field level, these systems support non-Abelian anyons, specifically Majorana zero modes localized at the ends of the chain. These quasiparticles, which are their own antiparticles, possess intriguing topological properties that can be used to encode and process quantum information Signature of Majorana modes have been observed in chains of magnetic atoms placed on top of superconductors [168, 215, 284]; however, there is no consensus on the topological origin of the observed boundary modes [134]. Following the work by Soldini et al. [228], where 2D regular lattices of magnetic impurities were created on a superconductor, we envision that this technique could be extended to construct quasicrystal and fractal lattices of magnetic impurities to realize topological superconductors in non-standard geometries. Quasiperiodic models, and especially the AA model, have attracted a lot of attention in the context of topological superconductivity [40, 64, 65, 288, 263, 281, 82]. The quasiperiodic potential provides an excellent testing ground for the robustness of the topological phase, as it does not break the protecting symmetries. Both analytic and numerical studies have explicitly confirmed that the topological phase and the Majorana quasiparticle persist in the presence of a weak quasiperiodic potential, but a strong incommensurate modulation drives a localization transition that destroys the topological phase. Notably, in some cases, quasiperiodic modulation of the chemical potential [82] or of the hopping [263] might even stabilize the topological phase.
@@ III.3 Cosmological models
The engineering of systems with curved geometries provides a new tool for studying gravitational physics and cosmology in the laboratory. As examples, we first consider the quantum simulation of gravitational effects in lattices which explicitly implement curvature through the appropriate design of connectivity. We then also mention the wide field of analog gravity: The analogy between sound propagation on a background hydrodynamic flow and field propagation in curved spacetime, first pointed out by Unruh [250], allows for studying gravitational phenomena even in flat systems. A similar analogy can also be made between light propagation in inhomogeneous media and field propagation in curved spacetime.
@@
The exceptional tunability of certain quantum systems, notably cold atomic systems and quantum fluid of light (i.e. non-linear transport of polariton BEC), renders them ideal for simulating particle flows that mimic curved spacetime for the system’s phononic excitations. For instance, the concept of “sonic black holes” within a BEC was proposed [90, 15, 46], leading to the realization of a polaritonic black hole [176] and the subsequent atomic realization of sonic black holes as experimental tools for investigating Hawking radiation [232, 177, 128]. The main idea behind these realizations is similar to the Unruh effect in classical fluids but now applied to quantum fluids: the “sonic horizon” separates a subsonic upstream region () from a supersonic downstream region (), as illustrated in Figs. 11(c,d), where is the speed of the BEC flow and are the speed of sound in the upstream and downstream regions, respectively. To probe the Hawking radiation in the case of 1D black holes, in Balbinot et al. [15], Carusotto et al. [47] has been proposed to use the second-order correlation function of density fluctuation, given by:
@@ IV Conclusion and Outlook
This colloquium has aimed at reviewing the state-of-the-art quantum engineering techniques and their reachable physical phenomena from a context that concentrates on exotic geometries. To this aim, we have covered various systems (atomic, electronic, and photonic ones), highlighting unique opportunities offered by different platforms and introducing general concepts such as Floquet engineering and synthetic dimensions. The second part of this colloquium has discussed specific implications of the exotic geometries on physical system behavior, including (i) localization phenomena, (ii) topological phenomena, (iii) analog black holes, and cosmology. These aspects already contain a rich variety of physical behavior on the single-body level, such as mobility edges in low dimensions (smaller than 3) or integer quantum Hall effect in high dimensions (larger than 3). While some aspects, such as topological classification in non-periodic crystals, remain an outstanding challenge, many features related to single-particle physics in exotic lattices have been both theoretically understood and experimentally realized. Much less explored, though, is the complicated subject of quantum many-body physics in such geometries. In this regard, the present colloquium article has provided a brief perspective on many-body localization. Theoretical research on this topic has indeed been ignited by the possibility of experimentally building synthetic quasicrystals that show this intriguing behavior or similar behavior (such as glassy phases). Yet there is open debate on the true existence of MBL phase, in particular in more than one dimension. Very little is known when it comes to topological many-body phenomena in quasicrystals or fractals, but the rich single-particle physics triggers hope for an even richer many-body scenario. Given the complexities in simulating such systems classically, the quantum simulators may develop their full potential of opening the door to deep questions of quantum many-body physics. With the engineering of exotic kinetic terms being achieved now, the basis has been set for future explorations of the interacting world of fractals, quasicrystals, and curved and higher-dimensional spaces.
@@ Acknowledgements.
T.G. and D.B. acknowledge fruitful discussions with Geza Giedke. D.B. acknowledges the support from the Spanish MICINN-AEI through Project No. PID2020-120614GB-I00 (ENACT), the Transnational Common Laboratory , and the Department of Education of the Basque Government through the project PIBA_2023_1_0007 (STRAINER). T.G. and D.B. acknowledge the financial support received from the IKUR Strategy under the collaboration agreement between the Ikerbasque Foundation and DIPC on behalf of the Department of Education of the Basque Government, and by the Gipuzkoa Provincial Council within the QUAN-000021-01 project. T.G. acknowledges funding by the Department of Education of the Basque Government through the project PIBA_2023_1_0021 (TENINT), by the Agencia Estatal de Investigación (AEI) through Proyectos de Generación de Conocimiento PID2022-142308NA-I00 (EXQUSMI), and that this work has been produced with the support of a 2023 Leonardo Grant for Researchers in Physics, BBVA Foundation. The BBVA Foundation is not responsible for the opinions, comments, and contents included in the project and/or the results derived therefrom, which are the total and absolute responsibility of the authors. H.S.N. is funded by the French National Research Agency (ANR) under the project README (ANR-22-CE09-0036-01). The work of C.W. is funded by the Cluster of Excellence “CUI: Advanced Imaging of Matter” of the Deutsche Forschungsgemeinschaft (DFG)—EXC 2056 —Project ID No. 390715994 and by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 802701. U.B. is also grateful for the financial support of the IBM Quantum Researcher Program.References
@@
- Li et al. [2023] Li, Meng, Chu Li, Linyu Yan, Qiang Li, Qihuang Gong, and Yan Li (2023), “Fractal photonic anomalous Floquet topological insulators to generate multiple quantum chiral edge states,” Light Sci. Appl. 12 (1), 262.

### v2

@@ III.2.1 Charge-pumping, Chern numbers, and gap labelling in 1D quasicrystals
In the case of 1D quasicrystals, the physical relevance of non-zero Chern numbers is evidenced by pumping experiments, where a quantized amount of “charge” is pumped through the chain when the cyclic Hamiltonian parameter is adiabatically changed. The amount of the pumped “charge” is defined by the Chern numbers of the filled bands. Such topological charge pumping has already been proposed by Thouless [262] in the context of 1D periodic (super)lattices. It has been realized in different photonic quasicrystals built from coupled waveguides in Kraus et al. [144], Verbin et al. [275] by adiabatically sweeping the phase of the quasiperiodic potential along the propagation axis. This setup has also served to visualize the presence or absence of localized states at the boundary between different quasicrystals [275], depending on whether the quasicrystals are topologically equivalent or not, see Fig. 10a. With cold atoms, one can fill the lowest band either as a fermionic band insulator [185] or a bosonic Mott insulator [164] and obtain a quantized charge transport of one site per modulation cycle.
@@ III.2.4 Topological edge states in fractals and quasicrystals
New topological phases are also expected or already observed in fractal and hyperbolic lattices. Whereas for standard geometries, the bulk-edge correspondence establishes an important relation between topological bulk behavior (Chern numbers) and topological edge states, extended bulk regions are lacking in fractal systems. Non-interacting topological systems (i.e., topological insulators and superconductors) on Sierpiński fractals have been studied theoretically in Brzezińska et al. [189], Pai and Prem [199], Iliasov et al. [122], Fremling et al. [94] , and in the Penrose tiling [86]. In these systems, the Hall conductivity is neither quantized nor proportional to the Chern number of the filled bands. However, splitting into various sub-bands makes it difficult to discern whether a band is filled. The fate of edge states is particularly intriguing since, on a fractal lattice, most sites are near some edge of the system. Topological edge states in a Sierpiński gasket have recently been observed in a photonic setup [31, 157], see Fig. 10(b).
@@
Theoretically simpler than the FQH effect are topological superconductors, such as the paradigmatic Kitaev chain [135]. Even on a mean-field level, these systems support non-Abelian anyons, specifically Majorana zero modes localized at the ends of the chain. These quasiparticles, which are their own antiparticles, possess intriguing topological properties that can be used to encode and process quantum information. Signature of Majorana modes have been observed in chains of magnetic atoms placed on top of superconductors [184, 231, 307]; however, there is no consensus on the topological origin of the observed boundary modes [147]. Following the work by Soldini et al. [245], where 2D regular lattices of magnetic impurities were created on a superconductor, we envision that this technique could be extended to construct quasicrystal and fractal lattices of magnetic impurities to realize topological superconductors in non-standard geometries. Quasiperiodic models, and especially the AA model, have attracted a lot of attention in the context of topological superconductivity [41, 68, 69, 311, 284, 303, 90]. The quasiperiodic potential provides an excellent testing ground for the robustness of the topological phase, as it does not break the protecting symmetries. Both analytic and numerical studies have explicitly confirmed that the topological phase and the Majorana quasiparticle persist in the presence of a weak quasiperiodic potential, but a strong incommensurate modulation drives a localization transition that destroys the topological phase. Notably, in some cases, quasiperiodic modulation of the chemical potential [90] or of the hopping [284] might even stabilize the topological phase.
@@ III.3 Cosmological models
The engineering of systems with curved geometries provides a new tool for studying gravitational physics and cosmology in the laboratory. As examples, we first consider the quantum simulation of gravitational effects in lattices, which explicitly implement curvature through the appropriate connectivity design. We then also mention the wide field of analog gravity: The analogy between sound propagation on a background hydrodynamic flow and field propagation in curved spacetime, first pointed out by Unruh [270], allows for studying gravitational phenomena even in flat systems. A similar analogy can also be made between light propagation in inhomogeneous media and field propagation in curved spacetime [134].
@@
The exceptional tunability of certain quantum systems, notably cold atomic systems and quantum fluid of light (i.e., non-linear transport of polariton BEC), renders them ideal for simulating particle flows that mimic curved spacetime for the system’s phononic excitations. For instance, the concept of “sonic black holes” within a BEC was proposed [98, 16, 48], leading to the realization of a polaritonic black hole [192] and the subsequent atomic realization of sonic black holes as experimental tools for investigating Hawking radiation [250, 193, 140]. The main idea behind these realizations is similar to the Unruh effect in classical fluids but now applied to quantum fluids: the “sonic horizon” separates a subsonic upstream region () from a supersonic downstream region (), as illustrated in Figs. 11(c,d), where is the speed of the BEC flow and are the speed of sound in the upstream and downstream regions, respectively. To probe the Hawking radiation in the case of 1D black holes, in Balbinot et al. [16], Carusotto et al. [49] has been proposed to use the second-order correlation function of density fluctuation, given by:
@@ IV Conclusion and Outlook
This colloquium has aimed at reviewing the state-of-the-art quantum engineering techniques and their reachable physical phenomena from a context that concentrates on exotic geometries. To this aim, we have covered various systems (atomic, electronic, and photonic ones), highlighting unique opportunities offered by different platforms and introducing general concepts such as Floquet engineering and synthetic dimensions. Many of these developments involve quantum control on the microscopic level and, therefore, resonate with strong efforts of building programmable quantum simulators. Such devices may then serve computational purposes and eventually demonstrate a quantum computational advantage, e.g. via boson sampling using a network of photons [316], or via quantum optimization using an Rydberg atoms in optical tweezer arrays [76].
In this colloquium, we have concentrated on specific physical implications of the exotic geometries. This includes (i) localization phenomena, (ii) topological phenomena, (iii) analog black holes, and cosmology. These aspects already contain a rich variety of physical behavior on the single-body level, such as mobility edges in low dimensions (smaller than 3) or integer quantum Hall effect in high dimensions (larger than 3). While some aspects, such as topological classification in non-periodic crystals, remain an outstanding challenge, many features related to single-particle physics in exotic lattices have been theoretically understood and experimentally realized. Much less explored, though, is the complicated subject of quantum many-body physics in such geometries. In this regard, the present colloquium article has provided a brief perspective on many-body localization. Theoretical research on this topic has indeed been ignited by the possibility of experimentally building synthetic quasicrystals that show this intriguing behavior or similar behavior (such as glassy phases). Yet there is open debate on the true existence of MBL phase, particularly in more than one dimension. Very little is known regarding topological many-body phenomena in quasicrystals or fractals. Still, the rich single-particle physics triggers hope for an even richer many-body scenario. Given the complexities in simulating such systems classically, the quantum simulators may develop their full potential of opening the door to deep questions of quantum many-body physics. With the engineering of exotic kinetic terms being achieved now, the basis has been set for future explorations of the interacting world of fractals, quasicrystals, and curved and higher-dimensional spaces.
@@ Acknowledgements.
[…]
@@
- Ahn et al. [2018] Ahn, Sung Joon, Pilkyung Moon, Tae-Hoon Kim, Hyun-Woo Kim, Ha-Chul Shin, Eun Hye Kim, Hyun Woo Cha, Se-Jong Kahng, Philip Kim, Mikito Koshino, Young-Woo Son, Cheol-Woong Yang, and Joung Real Ahn (2018), “Dirac electrons in a dodecagonal graphene quasicrystal,” Science 361 (6404), 782–786.
@@
- Cao et al. [2018] Cao, Yuan, Valla Fatemi, Shiang Fang, Kenji Watanabe, Takashi Taniguchi, Efthimios Kaxiras, and Pablo Jarillo-Herrero (2018), “Unconventional superconductivity in magic-angle graphene superlattices,” Nature 556 (7699), 43–50.
@@
- Chen et al. [2024] Chen, Anffany, Joseph Maciejko, and Igor Boettcher (2024), “Anderson localization transition in disordered hyperbolic lattices,” Phys. Rev. Lett. 133, 066101.
@@
- Darázs et al. [2014] Darázs, Zoltán, Anastasiia Anishchenko, Tamás Kiss, Alexander Blumen, and Oliver Mülken (2014), “Transport properties of continuous-time quantum walks on sierpinski fractals,” Phys. Rev. E 90, 032113.
@@
- Domany et al. [1983] Domany, Eytan, Shlomo Alexander, David Bensimon, and Leo P. Kadanoff (1983), “Solutions to the schrödinger equation on some fractal lattices,” Phys. Rev. B 28, 3110–3123.
@@
- Ebadi et al. [2022] Ebadi, S, A. Keesling, M. Cain, T. T. Wang, H. Levine, D. Bluvstein, G. Semeghini, A. Omran, J.-G. Liu, R. Samajdar, X.-Z. Luo, B. Nash, X. Gao, B. Barak, E. Farhi, S. Sachdev, N. Gemelke, L. Zhou, S. Choi, H. Pichler, S.-T. Wang, M. Greiner, V. Vuletić, and M. D. Lukin (2022), “Quantum optimization of maximum independent set using Rydberg atom arrays,” Science 376 (6598), 1209–1215.
@@
- Grossi e Fonseca et al. [2023] Grossi e Fonseca, André, Thomas Christensen, John D. Joannopoulos, and Marin Soljačić (2023), “Quasicrystalline weyl points and dense fermi-bragg arcs,” Phys. Rev. B 108, L121109.
- Fonseca et al. [2024] Fonseca, André Grossi, Sachin Vaidya, Thomas Christensen, Mikael C. Rechtsman, Taylor L. Hughes, and Marin Soljačić (2024), “Weyl points on nonorientable manifolds,” Phys. Rev. Lett. 132, 266601.
@@
- Gautier et al. [2021] Gautier, Ronan, Hepeng Yao, and Laurent Sanchez-Palencia (2021), “Strongly interacting bosons in a two-dimensional quasicrystal lattice,” Phys. Rev. Lett. 126, 110401.
@@
- Hao et al. [2024] Hao, Chen-Yue, Zhen Zhan, Pierre A. Pantaleón, Jia-Qi He, Ya-Xin Zhao, Kenji Watanabe, Takashi Taniguchi, Francisco Guinea, and Lin He (2024), “Robust flat bands in twisted trilayer graphene moiré quasicrystals,” Nat. Commun. 15 (1), 8437.
@@
- Kennes et al. [2021] Kennes, Dante M, Martin Claassen, Lede Xian, Antoine Georges, Andrew J. Millis, James Hone, Cory R. Dean, D. N. Basov, Abhay N. Pasupathy, and Angel Rubio (2021), “Moiré heterostructures as a condensed-matter quantum simulator,” Nature Physics 17 (2), 155–163.
@@
- Khveshchenko [2015] Khveshchenko, D V (2015), “Analogue holographic correspondence in optical metamaterials,” Europhysics Letters 109 (6), 61001.
@@
- Kosior and Sacha [2017] Kosior, Arkadiusz, and Krzysztof Sacha (2017), “Localization in random fractal lattices,” Phys. Rev. B 95, 104206.
@@
- Li et al. [2023a] Li, Hao, Yong-Yi Wang, Yun-Hao Shi, Kaixuan Huang, Xiaohui Song, Gui-Han Liang, Zheng-Yang Mei, Bozhen Zhou, He Zhang, Jia-Chi Zhang, Shu Chen, S. P. Zhao, Ye Tian, Zhan-Ying Yang, Zhongcheng Xiang, Kai Xu, Dongning Zheng, and Heng Fan (2023a), “Observation of critical phase transition in a generalized aubry-andré-harper model with superconducting circuits,” npj Quantum Information 9 (1), 40.
- Li et al. [2023b] Li, Meng, Chu Li, Linyu Yan, Qiang Li, Qihuang Gong, and Yan Li (2023b), “Fractal photonic anomalous Floquet topological insulators to generate multiple quantum chiral edge states,” Light Sci. Appl. 12 (1), 262.
@@
- Li et al. [2024] Li, Yanxing, Fan Zhang, Viet-Anh Ha, Yu-Chuan Lin, Chengye Dong, Qiang Gao, Zhida Liu, Xiaohui Liu, Sae Hee Ryu, Hyunsue Kim, Chris Jozwiak, Aaron Bostwick, Kenji Watanabe, Takashi Taniguchi, Bishoy Kousa, Xiaoqin Li, Eli Rotenberg, Eslam Khalaf, Joshua A. Robinson, Feliciano Giustino, and Chih-Kang Shih (2024), “Tuning commensurability in twisted van der Waals bilayers,” Nature 625 (7995), 494–499.
@@
- Lundblad et al. [2023] Lundblad, Nathan, David C Aveline, Antun Balaž, Elliot Bentine, Nicholas P Bigelow, Patrick Boegel, Maxim A Efremov, Naceur Gaaloul, Matthias Meister, Maxim Olshanii, Carlos A R Sá de Melo, Andrea Tononi, Smitha Vishveshwara, Angela C White, Alexander Wolf, and Barry M Garraway (2023), “Perspective on quantum bubbles in microgravity,” Quantum Science and Technology 8 (2), 024003.
@@
- Sierant et al. [2024] Sierant, Piotr, Maciej Lewenstein, Antonello Scardicchio, Lev Vidmar, and Jakub Zakrzewski (2024), “Many-body localization in the age of classical computing,” arXiv:2403.07111 [cond-mat.dis-nn] .
@@
- Stampfli [1986] Stampfli, P (1986), “A dodecagonal quasiperiodic lattice in two dimensions.” Helv. Phys. Acta 59, 1260–1263.
@@
- Tarquini et al. [2017] Tarquini, E, G. Biroli, and M. Tarzia (2017), “Critical properties of the Anderson localization transition and the high-dimensional limit,” Phys. Rev. B 95, 094204.
@@
- Tononi et al. [2022] Tononi, Andrea, Axel Pelster, and Luca Salasnich (2022), “Topological superfluid transition in bubble-trapped condensates,” Phys. Rev. Res. 4, 013122.
@@
- Uri et al. [2023] Uri, Aviram, Sergio C. de la Barrera, Mallika T. Randeria, Daniel Rodan-Legrain, Trithep Devakul, Philip J. D. Crowley, Nisarga Paul, Kenji Watanabe, Takashi Taniguchi, Ron Lifshitz, Liang Fu, Raymond C. Ashoori, and Pablo Jarillo-Herrero (2023), “Superconductivity and strong interactions in a tunable moiré quasicrystal,” Nature 620 (7975), 762–767.
@@
- Wang [1995] Wang, Xiang Rong (1995), “Localization in fractal spaces: Exact results on the Sierpinski gasket,” Phys. Rev. B 51, 9310–9313.
@@
- Yao et al. [2018] Yao, Wei, Eryin Wang, Changhua Bao, Yiou Zhang, Kenan Zhang, Kejie Bao, Chun Kai Chan, Chaoyu Chen, Jose Avila, Maria C. Asensio, Junyi Zhu, and Shuyun Zhou (2018), “Quasicrystalline 30° twisted bilayer graphene as an incommensurate superlattice with strong interlayer coupling,” Proc. Natl. Acad. Sci. U.S.A. 115 (27), 6928–6933.
@@
- Zhang et al. [2021] Zhang, Ren, Chenwei Lv, Yangqian Yan, and Qi Zhou (2021), “Efimov-like states and quantum funneling effects on synthetic hyperbolic surfaces,” Science Bulletin 66 (19), 1967–1972.
@@
- Zhong et al. [2020] Zhong, Han-Sen, Hui Wang, Yu-Hao Deng, Ming-Cheng Chen, Li-Chao Peng, Yi-Han Luo, Jian Qin, Dian Wu, Xing Ding, Yi Hu, Peng Hu, Xiao-Yan Yang, Wei-Jun Zhang, Hao Li, Yuxuan Li, Xiao Jiang, Lin Gan, Guangwen Yang, Lixing You, Zhen Wang, Li Li, Nai-Le Liu, Chao-Yang Lu, and Jian-Wei Pan (2020), “Quantum computational advantage using photons,” Science 370 (6523), 1460–1463.
@@
- Zhu et al. [2024] Zhu, Zhaoxuan, Shengjie Yu, Dean Johnstone, and Laurent Sanchez-Palencia (2024), “Localization and spectral structure in two-dimensional quasicrystal potentials,” Phys. Rev. A 109, 013314.